OpenStudy (anonymous):

solve differential equation using exact method. (x+y)^2dx+(2xy+x^2-3)dy=0 for the intial condition of y(1)=1

OpenStudy (anonymous):

my answer was xy^2+x^2y-3y+x^3/3

OpenStudy (anonymous):

=C

OpenStudy (anonymous):

am i right ?

OpenStudy (anonymous):

not sure how to apply initial condition

zepdrix (zepdrix):

Ummm yah those are the same terms I'm coming up with! \[\Large\int\limits (x+y)^2dx+\int\limits (2xy+x^2-3)dy=\int\limits 0\]Integrating each side gives us,\[\Large\sf \frac{1}{3}x^3+xy^2+x^2y-3y=c\]y(1)=1, you can interpret this as, x=1 when y=1. So just plug in a bunch of 1's to solve for c,\[\Large\sf \frac{1}{3}+1+1-3=c\]

OpenStudy (anonymous):

i dont know what am smoking ....that so passed my mind

OpenStudy (anonymous):

dont do DFQ for 2 days n i forget everything

zepdrix (zepdrix):

heh XD

OpenStudy (anonymous):

thx u